• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dhamilton

    Log entry time 00:03:42 on April 6,2008

    Entry number 230221

    keyword=Swing Shift Summary

    Swing Shift Summary
    SL: David Hamilton
    TO: Jin Huang

    I took over from Richard around 17:30.

    17:30 - Problems with HV trips when BB current is off.

    18:15 - Move to pi0 kinematics (HRS 20.5 deg, 960 MeV/c, BB 390A) with BigBite sieve IN.
    Run 2351 (0.5M, LH2, BB 390A, BB Sieve IN)
    Run 2352 (0.5M, LH2, BB 390A, BB Sieve IN)

    18:50 - Move Hall status to controlled access, in order to remove the BigBite sieve slit.

    19:30 - Hall status changed back to beam permit, following removal of the BigBite sieve collimator.

    19:45 - target changed to Ta. HRSl momentum set to 1194 MeV/c (Q2 and Q3 cycled through 1400A). Raster off. Prescales changed for T3 and T4 running.
    Run 2356 (0.25M, Ta, T3 and T4, replay ok)

    20:20 - target changed to C.
    Run 2357 (0.5M, C, T3 and T4, replay ok)

    20:50 - target changed to 6cm LH2, HRSl central momentum set to 1104.9 MeV/c.
    Run 2358 (CODA rate too high - junk run)
    Run 2359 (CODA rate too high - junk run)
    Run 2360 (0.5M, LH2, T3 and T4, replay ok)
    Run 2361 (0.5M, LH2, T3 and T4, replay ok)

    21:20 - move to production kinematics.
    HRSl angle = 20.5 degrees
    HRSl pcent = 960 MeV/c
    BB angle = 54 degress
    BB coil I = 390 A
    Target = 6cm LH2
    Raster off
    5 microamps
    Run 2362 (0.5M, LH2, production, replay ok)
    Run 2363 (0.5M, LH2, production, replay ok)

    22:25 - HRS Right Arm Q2 power supply trip. RC advises continuing production running. The power supply can be reset locally during the next controlled access or during the day shift.

    Run 2364 (0.5M, LH2, production, replay ok)
    Run 2365 (0.1M, LH2, production, replay ok) ended early due to HV trip

    23:00 - BB HV trip fixed. Resume data-taking.
    Run 2366 (0.5M, LH2, production, replay ok)
    Run 2367 (0.5M, LH2, production, replay ok)
    Run 2368 (0.5M, LH2, production, replay ok)

    23:45 - BB HV GUI crashed. When we brought it back up, we found that BB MWDC HV had tripped. Runs 2369 and 2370 are junk. We confirmed this in the replay.

    Run 2371 continues through the end of the shift.